# Unsupervised Visual Chain-of-Thought Reasoning via Preference Optimization

### v1

@@ Abstract
Chain-of-thought (CoT) reasoning greatly improves the interpretability and problem-solving abilities of multimodal large language models (MLLMs). However, existing approaches are focused on text CoT, limiting their ability to leverage visual cues. Visual CoT remains underexplored, and the only work [28] is based on supervised fine-tuning (SFT) that relies on extensive labeled bounding-box data and is hard to generalize to unseen cases. In this paper, we introduce Unsupervised Visual CoT (UV-CoT), a novel framework for image-level CoT reasoning via preference optimization. UV-CoT performs preference comparisons between model-generated bounding boxes (one is preferred and the other is dis-preferred), eliminating the need for bounding-box annotations. We get such preference data by introducing an automatic data generation pipeline. Given an image, our target MLLM (e.g., LLaVA-1.5-7B) generates seed bounding boxes using a template prompt and then answers the question using each bounded region as input. An evaluator MLLM (e.g., OmniLLM-12B) ranks the responses, and these rankings serve as supervision to train the target MLLM with UV-CoT by minimizing negative log-likelihood losses. By emulating human perception–identifying key regions and reasoning based on them–UV-CoT can improve visual comprehension, particularly in spatial reasoning tasks where textual descriptions alone fall short. Our experiments on six datasets demonstrate the superiority of UV-CoT, compared to the state-of-the-art textual and visual CoT methods. Our zero-shot testing on four unseen datasets shows the strong generalization of UV-CoT. The code is available in https://kesenzhao.github.io/my_project/projects/UV-CoT.html.
@@ 1 Introduction
With the recent advancements in multimodal large language models (MLLMs) [17, 18, 3, 39], many efforts have been made to incorporate text CoT reasoning [34, 13, 42, 8] to handle complex vision-language tasks [43, 37, 36]. However, the visual understanding ability of MLLM is limited by fixed-granularity image processing, i.e., the MLLM can’t dynamically adjust focus across different spatial regions of the input image, even when guided by text CoT prompts [9]. This underscores the critical need to explicitly integrate visual cues into the CoT process.
Figure 1: Comparison of Visual CoT [28] and our UV-CoT. Left: Visual CoT relies on human-annotated bounding boxes to identify key regions. The model is trained via supervised fine-tuning to maximize the likelihood of the labeled data. Right: UV-CoT eliminates the need for human annotation. Given an image, the target model generates seed bounding boxes and answers questions based on these regions, respectively. An evaluator MLLM then scores the responses as a proxy for assessing region quality. Lastly, the target model is optimized via preference optimization by maximizing the likelihood of preferred regions over dispreferred ones.
A very recent work called Visual-CoT [28] has made an initial attempt towards this goal. The model is trained using supervised fine-tuning (SFT) with human-labeled bounding boxes that indicate the key image regions relevant to the question. Following this, it performs its multimodal CoT approach with human-annotated reasoning steps by associating textual inputs with the detected regions. An overview of Visual CoT is presented in Fig. 1. However, it has two key drawbacks: (1) it relies on large-scale, high-quality labeled data, making it costly and hard to scale; and (2) SFT learns only from positive examples (i.e., the labeled training data), limiting its ability to generalize to unseen or ambiguous scenarios where intermediate reasoning or dynamic interpretation is needed.
@@
Our method UV-CoT, as shown in Fig. 1, differs from Visual CoT [28] by adopting an unsupervised approach with contrastive preference data. We design an automatic two-step pipeline to generate this data: 1) Region Generation: Given an image, the target model generates multiple seed bounding boxes using a template prompt. It then answers the question by processing each bounded region together with the question as input. 2) Quality Assessment: An evaluator MLLM scores the responses, using these scores as proxies to measure the quality of the regions. Unlike traditional DPO, we propose Score-DPO (sDPO), which not only ranks preference data (i.e., preferred and dis-preferred responses shown in Fig. 1) but also assigns preference scores. This scoring enables more precise optimization based on score differences. During UV-CoT training, the rankings of the preference data act as supervision by minimizing negative log-likelihood loss, while the scores define the decision margin. By mimicking human perception–first identifying key regions, then reasoning over them–UV-CoT significantly improves visual comprehension, especially in spatial reasoning tasks where text-based methods fall short. By leveraging unsupervised data in a contrastive way, UV-CoT also shows strong generalization ability when tested on unseen datasets. We will show the evidence in the experimental sections.
Our contributions in this paper thus include: 1) an automatic pipeline for generating high-quality preference data, enabling robust and scalable preference learning of UV-CoT; 2) an improved version of DPO in UV-CoT by integrating the degree of preference for visual regions, allowing the model to distinguish key regions more precisely and improve its reasoning ability; and 3) extensive experiments on multiple challenging datasets, demonstrating state-of-the-art performance of UV-CoT on four benchmarks and strong generalization to four unseen test datasets.
@@ 2 Related Works
Chain-of-Thought in LLMs and MLLMs. LLMs [19, 32, 3] show strong inferential abilities, with CoT [34, 7] improving their performance by guiding models through intermediate reasoning steps. Both manually designed [34] and self-generated [13, 42] reasoning approaches have proven effective in improving LLM performance. In contrast, MLLMs struggle with reasoning due to differences in processing textual and visual data. Multimodal CoT methods [43, 37, 36] attempt to address this by transforming multimodal inputs into a unified textual format, enabling LLMs to perform CoT at the text level. However, this transformation introduces significant information loss and prevents the models from capturing key visual details [43]. For example, LLaVA-CoT [36] leverages GPT-4o to summarize questions and image captions but suffers from weak optical character recognition and sometimes hallucinations.
A very recent work, Visual CoT [28], improves the MLLM reasoning by introducing CoT methods at the image level. This approach involves scanning the entire image, identifying key references, and then focusing the model on specific regions for reasoning. Despite its improvements, Visual CoT is heavily based on costly human-annotated data. In contrast, our UV-CoT framework utilizes unsupervised preference optimization with auto-generated preference data, eliminating the need for manual annotations.
Preference Learning in LLMs and MLLMs. RLHF [31] aligns LLMs with human preferences by training a reward model via contrastive response evaluations [45]. To reduce reliance on human annotations, RLAIF [15] leverages pre-trained LLMs for preference label generation. However, RL-based fine-tuning faces stability and efficiency challenges. Direct Preference Optimization (DPO) [26] addresses this by directly linking reward functions to optimal policies, eliminating reward model fine-tuning. Further improvements include IPO [2], which mitigates overfitting with a bounded preference function, and KTO [6], which removes the need for paired preference data, relying instead on single examples labeled as either “good” or “bad”.
@@ 3.1 Preference Data Generation
Given a target model $f_{\mathsf{tar}}$, an evaluator model $f_{\mathsf{eval}}$, and an image-question pair $x$, we illustrate how to construct $n$ preference data points. The evaluator model $f_{\mathsf{eval}}$ is a slightly stronger MLLM than the target model $f_{\mathsf{tar}}$, providing a better assessment of the responses [^1] Please note that the target MLLM can also be used as the evaluator, validated in Tab. 5.. Assume we have $T$ steps in CoT reasoning, we generate preference data for $T$ times on the way, as described in Algorithm 1. At each timestep $t$ (i.e., a reasoning step $t$), the process includes four stages: Response Generation, Response Evaluation, Pair Construction, and Response Selection.
Response Generation. The goal of this stage is to generate seed bounding boxes and produce intermediate responses to the question using the target model. Here, a “response” refers to any model output at an intermediate step, not necessarily the final answer to the question. We denote the model’s response at timestep $t$ as $y_{t}$, with the initial input $x$ represented as $y_{0}$. To create diverse bounding boxes and subsequent responses, we apply stochastic decoding on the target model $f_{\mathsf{tar}}$ with $n$ different seeds, resulting in a set of responses ${y_{t}^{i}}_{i=1}^{n}$.
Response Evaluation. This stage evaluates the quality of all generated responses. The evaluator model not only scores each individual response but also counts how this response influences the quality of its next response in the chain. This cumulative evaluation approach helps quantify the impact of each bounded region on the overall reasoning process. Below, we elaborate on the formulation. At timestep $t$, the evaluator assign scores for $y_{t}^{i}$ as follows:
@@
where $s_{\mathsf{nxt}}^{i}$ reflects the impact on the next response and $\gamma>0$ is a hyperparameter to combine the scores of current and the next response, with $\gamma=0$ at the last step. We estimate the expectation $\mathbb{E}[\cdot]$ by randomly sampling next responses.
Pair Construction. At each timestep $t$, we randomly select $k$ (preferred and dis-preferred) pairs from $\{y_{t}^{i}\}_{i=1}^{n}$. For a single pair, we concatenate the preferred response with the past response chain $y_{0:t-1}$ (preserved after $t\!-\!1$ timesteps) and get a “preferred chain” denoted as $y_{t}^{w}$, and then we concatenate the dis-preferred response in the same way to get a “dis-preferred chain” denoted as $y_{t}^{l}$. The pair of chains also includes their respective scores, and they are represented as $\{y_{w},s_{w},y_{l},s_{l}\}$. The overall $k$ pairs of chains compose the preference dataset $\mathcal{D}_{t}$.
Response Selection. The abovementioned “past response chain $y_{0:t-1}$” is unique and is concatenated by the highest-scoring response at timestep $t\!-\!1$ and the preserved chain at timestep $t\!=\!2$, i.e., $y_{0:t-2}$. In other words, when finishing each timestep process, we preserve only the best chain and use it for the next step.
@@ 3.2 Unsupervised Learning of UV-CoT
$\mathcal{L}_{sDPO}(\theta)=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.$ (2)
@@
Using the Bradley-Terry [4] model, the standard DPO aims to minimize the negative log-likelihood of the difference of rewards:
@@
This result follows from the definition of Gumbel random variables [1] and the Gumbel-max trick [20]. We provide a detailed proof in the Appendix B. By maximizing the log-likelihood of this objective, we obtain our proposed loss function. The Gumbel distribution models the extreme values of a variable, while $\Delta_{r}$ quantifies the degree of difference between preference pairs. Thus, our loss function explicitly optimizes preference learning by distinguishing not only the order but also the magnitude of preference differences.
@@
MLLM DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393 LLaVA-1.5-13B 0.225 0.543 0.169 0.607 0.506 0.512 0.427 MiniCPM-o-8B 0.232 0.529 0.175 0.618 0.495 0.521 0.428 OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443 Visual-CoT-7B (100% label) 0.294 0.673 0.194 0.652 0.546 0.532 0.482 UV-CoT (0% label) 0.265 0.686 0.173 0.632 0.536 0.548 0.473 UV-CoT (10% label) 0.283 0.711 0.198 0.649 0.568 0.553 0.494
Table 1: Overall comparison of different models on six evaluation benchmarks. The best result is bold, the second-best is underlined. “%” indicates the percentage of supervised data used in Visual-CoT. By default, our UV-CoT uses only unsupervised data.
@@ 4 Experiments
MLLM DUDE SROIE Visual7w Average LLaVA-1.5-7B 0.165 0.147 0.340 0.217 LLaVA-1.5-13B 0.174 0.159 0.352 0.228 MiniCPM-o-8B 0.182 0.165 0.341 0.229 OmniLMM-12B 0.194 0.166 0.357 0.239 Visual-CoT-7B 0.206 0.181 0.397 0.261 UV-CoT 0.241 0.184 0.432 0.286 UV-CoT 0.253 0.227 0.455 0.312
Table 2: Zero-short experiments on DUDE, SROIE and Visual7w. The best result is bold, the second-best is underlined. ‘UV-CoT’ denotes our model trained with additional unlabeled preference data from the three zero-shot datasets.
MLLM Attributes GPT4V-hard OCR Average LLaVA-1.5-7B 0.317 0 0.1 0.139 LLaVA-1.5-13B 0.326 0.118 0.133 0.192 MiniCPM-o-8B 0.322 0.118 0.1 0.180 OmniLMM-12B 0.326 0.118 0.167 0.204 Visual-CoT-7B 0.330 0.118 0.593 0.347 UV-CoT 0.352 0.176 0.677 0.402
Table 3: High-resolution image reasoning on V Bench. The best result is bold.
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average UV-CoT (10% labels) 0.283 0.711 0.198 0.649 0.568 0.553 0.494 IF: w/o UV-CoT 0.149 0.574 0.160 0.585 0.509 0.522 0.417 IF: UV-CoT w/ G.T. BBox 0.528 0.769 0.504 0.655 0.664 0.585 0.618 Tr: w/ naive DPO 0.258 0.695 0.189 0.623 0.552 0.534 0.475 Tr: w/o iterative learning 0.256 0.671 0.162 0.609 0.523 0.531 0.459 Ge: w/o $\gamma$ 0.247 0.539 0.152 0.551 0.484 0.460 0.406

### v2

@@ Abstract
Chain-of-thought (CoT) reasoning greatly improves the interpretability and problem-solving abilities of multimodal large language models (MLLMs). However, existing approaches focus on text CoT, limiting their ability to leverage visual cues. Visual CoT remains underexplored, and the only work [35] is based on supervised fine-tuning that relies on extensive labeled bounding-box data and is hard to generalize to unseen cases. In this paper, we introduce Unsupervised Visual CoT (UV-CoT), a novel framework for image-level CoT reasoning via preference optimization. UV-CoT performs preference comparisons between model-generated bounding boxes (one is preferred and the other is dis-preferred), eliminating the need for bounding-box annotations. We get such preference data by introducing an automatic data generation pipeline. Given an image, our target MLLM (e.g., LLaVA-1.5-7B) generates seed bounding boxes using a template prompt and then answers the question using each bounded region as input. An evaluator MLLM (e.g., OmniLLM-12B) ranks the responses, and these rankings serve as supervision to train the target MLLM with UV-CoT by minimizing negative log-likelihood losses. By emulating human perception–identifying key regions and reasoning based on them–UV-CoT can improve visual comprehension, particularly in spatial reasoning tasks where textual descriptions alone fall short. Our experiments on six datasets demonstrate the superiority of UV-CoT, compared to the state-of-the-art textual and visual CoT methods. Our zero-shot testing on four unseen datasets shows the strong generalization of UV-CoT.
@@ 1 Introduction
With the recent advancements in multimodal large language models (MLLMs) [21, 22, 3, 45], many efforts have been made to incorporate text CoT reasoning [40, 17, 49, 9] to handle complex vision-language tasks [50, 43, 42]. However, the visual understanding ability of MLLM is limited by fixed-granularity image processing, i.e., the MLLM cannot dynamically adjust focus across different spatial regions of the input image, even when guided by text CoT prompts [11]. This underscores the critical need to explicitly integrate visual cues into the CoT process.
Figure 1: Comparison of Visual-CoT [35] and our UV-CoT. Left: Visual-CoT relies on human-annotated bounding boxes to identify key regions. The model is trained via supervised fine-tuning to maximize the likelihood of the labeled data. Right: UV-CoT eliminates the need for human annotation. Given an image, the target model generates seed bounding boxes and answers questions based on these regions, respectively. An evaluator MLLM then scores the responses as a proxy for assessing region quality. Lastly, the target model is optimized via preference optimization by maximizing the likelihood of preferred regions over dis-preferred ones.
A very recent work, Visual-CoT [35], has made an initial attempt towards this goal. The model is trained using supervised fine-tuning (SFT) with human-labeled bounding boxes that indicate the key image regions relevant to the question. It performs the multimodal CoT approach with human-annotated reasoning steps by associating textual inputs with the detected regions. An overview of Visual-CoT is presented in Fig. 1. However, it has two key drawbacks: (1) it relies on large-scale, high-quality labeled data, making it costly and hard to scale; and (2) SFT learns only from positive examples (i.e., the labeled data), limiting its ability to generalize to unseen or ambiguous scenarios where intermediate reasoning or dynamic interpretation is needed.
@@
Our method UV-CoT, as shown in Fig. 1, differs from Visual-CoT [35] by adopting an unsupervised approach with contrastive preference data. We design an automatic two-step pipeline to generate this data: 1) Region Generation: Given an image, the target model generates multiple seed bounding boxes using a template prompt. Then it answers the question by processing each bounded region together with the question as input. 2) Quality Assessment: An evaluator MLLM scores the responses, using these scores as proxies to measure the quality of the regions. Unlike traditional DPO, we propose Score-DPO (sDPO), which not only ranks preference data (i.e., preferred and dis-preferred responses shown in Fig. 1) but also assigns preference scores. This scoring enables more precise optimization based on score differences. During UV-CoT training, the rankings of the preference data act as supervision by minimizing negative log-likelihood loss, while the scores define the decision margin. By mimicking human perception–first identifying key regions, then reasoning over them–UV-CoT significantly improves visual comprehension, especially in spatial reasoning tasks where text-based methods fall short. By leveraging unsupervised data in a contrastive way, UV-CoT also shows strong generalization ability when tested on unseen datasets.
Our contributions in this paper include: 1) an automatic pipeline for generating high-quality preference data, enabling robust and scalable preference learning of UV-CoT; 2) an improved version of DPO by integrating the degree of preference for visual regions, allowing the model to distinguish key regions more precisely; and 3) extensive experiments on multiple challenging datasets, demonstrating state-of-the-art performance of UV-CoT on four benchmarks and strong generalization to four unseen test datasets.
@@ 2 Related Works
Chain-of-thought in LLMs and MLLMs. LLMs [24, 38, 3, 29, 10, 14] with CoT [40, 8, 5] show strong inferential abilities by introducing intermediate reasoning steps. Both manually designed [40] and self-generated [17, 49] reasoning approaches have proven effective. In contrast, MLLMs rely on image encoders [32, 54, 15, 52, 53] to extract visual features but often struggle with reasoning due to inherent differences in how textual and visual data are processed [51, 23, 48]. Multimodal CoT methods [50, 43, 42] attempt to address this by transforming multimodal inputs into a unified textual format, enabling LLMs to perform CoT at the text level. However, this transformation introduces significant information loss and prevents the models from capturing key visual details [50]. For example, LLaVA-CoT [42] leverages GPT-4o to summarize questions and image captions but suffers from weak optical character recognition and sometimes hallucinations.
A very recent work, Visual-CoT [35], improves the MLLM reasoning by introducing CoT methods at the image level. This approach involves scanning the entire image, identifying key references, and then focusing the model on specific regions for reasoning. Despite its improvements, Visual-CoT is heavily based on costly human-annotated data. In contrast, our UV-CoT framework utilizes unsupervised preference optimization with auto-generated preference data, eliminating the need for manual annotations.
Preference learning in LLMs and MLLMs. RLHF [56] aligns LLMs with human preferences by training a reward model via contrastive response evaluations. To reduce reliance on human annotations, RLAIF [19] leverages pretrained LLMs for preference label generation. However, RL-based fine-tuning faces stability and efficiency challenges. Direct Preference Optimization (DPO) [33] addresses this by directly linking reward functions to optimal policies, eliminating reward model fine-tuning. Further improvements include IPO [2], which mitigates overfitting with a bounded preference function, and KTO [7], which removes the need for paired preference data, relying instead on single examples labeled as either ‘good’ or ‘bad’.
@@ 3.1 Preference Data Generation
Given a target model $f_{\mathsf{tar}}$, an evaluator model $f_{\mathsf{eval}}$ (the target MLLM can also serve as the evaluator, validated in Tab. 5), and an image-question pair $x$, we illustrate how to construct $n$ preference data points. Assuming there are $T$ steps in CoT reasoning process, we generate preference data for $T$ times on the way, as described in Algorithm 1. At each timestep $t$ (i.e., a reasoning step $t$), the process includes four stages: Response Generation, Response Evaluation, Pair Construction, and Response Selection.
Response generation. The goal of this stage is to generate seed bounding boxes and produce intermediate responses to the question using the target model. Here, a ‘response’ refers to any model output at an intermediate step, not necessarily the final answer to the question. We denote the model’s response at timestep $t$ as $y_{t}$, with the initial input $x$ represented as $y_{0}$. To encourage diversity in the bounding boxes and subsequent responses, we apply stochastic decoding to the target model $f_{\mathsf{tar}}$ with $n$ different random seeds, resulting in a set of responses $\{y_{t}^{i}\}_{i=1}^{n}$.
Response evaluation. This stage evaluates the quality of all generated responses. The evaluator model not only scores each individual response but also counts how this response influences the quality of its next response in the chain. This cumulative evaluation approach helps quantify the impact of each bounded region on the overall reasoning process. Below, we elaborate on the formulation. At timestep $t$, the evaluator assigns scores for $y_{t}^{i}$ as follows:
@@
where $s_{\mathsf{nxt}}^{i}$ reflects the impact on the next response and $\gamma>0$ is a hyperparameter to combine the current and next response scores, with $\gamma=0$ at the last step. We estimate the expectation $\mathbb{E}[\cdot]$ by randomly sampling next responses.
Pair construction. At each timestep $t$, we randomly select $k$ (preferred and dis-preferred) pairs from $\{y_{t}^{i}\}_{i=1}^{n}$. For a single pair, we concatenate the preferred response with the past response chain $y_{0:t-1}$ (preserved after $t\!-\!1$ timesteps) and get a “preferred chain” denoted as $y_{t}^{w}$, and then we concatenate the dis-preferred response in the same way to get a ‘dis-preferred chain’ denoted as $y_{t}^{l}$. The pair of chains also includes their respective scores, and they are represented as $\{y_{w},s_{w},y_{l},s_{l}\}$. The overall $k$ pairs of chains compose the preference dataset $\mathcal{D}_{t}$.
Response selection. The abovementioned ‘past response chain $y_{0:t-1}$’ is unique and is concatenated by the highest-scoring response at timestep $t\!-\!1$ and the preserved chain at timestep $t\!=\!2$, i.e., $y_{0:t-2}$. In other words, when finishing each timestep process, we preserve only the best chain and use it for the next step.
@@ 3.2 Unsupervised Learning of UV-CoT
$\mathcal{L}_{\text{sDPO}}(\theta)=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.$ (2)
@@
Using the Bradley-Terry [4] model, the standard DPO aims to minimize the negative log-likelihood of the difference of rewards between paired responses:
@@
This result follows from the definition of Gumbel random variables [1] and the Gumbel-max trick [25]. We provide a detailed proof in Appendix B. By maximizing the log-likelihood, we obtain our proposed loss function. The Gumbel distribution models the extreme values of a variable, while $\Delta_{r}$ quantifies the degree of difference between preference pairs. Thus, our loss function explicitly optimizes preference learning by distinguishing not only the order but also the magnitude of preference differences.
@@
MLLM DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average
LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393
LLaVA-1.5-13B 0.225 0.543 0.169 0.607 0.506 0.512 0.427
MiniCPM-o-8B 0.232 0.529 0.175 0.618 0.495 0.521 0.428
OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443
Visual-CoT-7B (100% label) 0.294 0.673 0.194 0.652 0.546 0.532 0.482
UV-CoT (0% label) 0.265 0.686 0.173 0.632 0.536 0.548 0.473
UV-CoT (10% label) 0.283 0.711 0.198 0.649 0.568 0.553 0.494
Table 1: Overall comparison of different models on six evaluation benchmarks. The best result is bold, the second-best is underlined. ‘%’ indicates the percentage of supervised data used in Visual-CoT. By default, our UV-CoT uses only unsupervised data.
@@ 4 Experiments
MLLM DUDE SROIE Visual7w Average
LLaVA-1.5-7B 0.165 0.147 0.340 0.217
LLaVA-1.5-13B 0.174 0.159 0.352 0.228
MiniCPM-o-8B 0.182 0.165 0.341 0.229
OmniLMM-12B 0.194 0.166 0.357 0.239
Visual-CoT-7B 0.206 0.181 0.397 0.261
UV-CoT 0.241 0.184 0.432 0.286
UV-CoT 0.253 0.227 0.455 0.312
Table 2: Zero-shot experiments on DUDE, SROIE and Visual7w. ‘UV-CoT’ denotes our model trained with additional unlabeled preference data from the three zero-shot datasets.
4.1 Setup
Datasets. For a comprehensive evaluation, we adopt ten datasets spanning five domains: (1) Text/Document: DocVQA [26], TextVQA [36], DUDE [39], and SROIE [12]. (2) Chart: InfographicsVQA [27]. (3) General VQA: Flickr30k [31] and Visual7W [55]. (4) Relation Reasoning: VSR [20] and GQA [13]. (5) High-Resolution Image Reasoning: V Bench [41]. Notably, we also provide a model trained on data excluding DUDE, SROIE, Visual7W, and V Bench, which is used to evaluate zero-shot performance. See Appendix C.1 for details.
Evaluation. Following prior work [22], we prompt GPT-4o [28] to assign a score between 0 and 1, with higher scores indicating better prediction. See details in Appendix C.2.
Baselines. We compare UV-CoT with five baselines. LLaVA-1.5-{7B, 13B} [21] and OmniLMM-12B are strong general baselines. MiniCPM-o-8B [44] adopts adaptive visual encoding for fine-grained understanding and Visual-CoT-7B [35] learns image-level CoT via SFT.
Implementation details. We use LLaVA-1.5-7B as the target model and OmniLMM-12B as the evaluator. To ensure scalability, we avoid using GPT-4 as the evaluator, preventing constraints imposed by high API costs. We implement iterative learning of UV-CoT over four iterations, utilizing a total of 249K preference data pairs. Notably, UV-CoT achieves higher data efficiency than Visual-CoT, which uses 376K data pairs. For each iteration, we train the model with AdamW optimizer for 4 epochs with a learning rate of $0.15\times 10^{-7},\beta=$ 0.1, and a batch size of 8. In total, data generation takes 80 hours and training requires 60 hours, both conducted on an 8$\times$×A100 40GB machine. Additionally, we provide a variant UV-CoT trained with extra SFT on 10% of the labeled Visual-CoT data, denoted as UV-CoT (10%).
MLLM Attributes GPT4V-hard OCR Average
LLaVA-1.5-7B 0.317 0 0.1 0.139
LLaVA-1.5-13B 0.326 0.118 0.133 0.192
MiniCPM-o-8B 0.322 0.118 0.1 0.180
OmniLMM-12B 0.326 0.118 0.167 0.204
Visual-CoT-7B 0.330 0.118 0.593 0.347
UV-CoT 0.352 0.176 0.677 0.402
Table 3: Zero-shot experiments on V Bench (high-resolution image reasoning task, average resolution $2246\times 1582$).
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average
UV-CoT (10% labels) 0.283 0.711 0.198 0.649 0.568 0.553 0.494
IF: w/o UV-CoT 0.149 0.574 0.160 0.585 0.509 0.522 0.417
IF: UV-CoT w/ G.T. BBox 0.528 0.769 0.504 0.655 0.664 0.585 0.618
Tr: w/ naive DPO 0.258 0.695 0.189 0.623 0.552 0.534 0.475
Tr: w/o iterative learning 0.256 0.671 0.162 0.609 0.523 0.531 0.459
Ge: w/o $\gamma$ 0.247 0.539 0.152 0.551 0.484 0.460 0.406